

********************************************************************
UARS WEEKLY STATUS REPORT - 18 Jan 2002
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods.
***********
SUSIM
***********
SUSIM completed daily mid resolution and MgII solar scans during
periods of predicted UARS telemetry. We are awaiting the corrected
level 0 data from the PACOR-A system.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM continues to receive NMC data files from NOAA. No problems have
been encountered.
An initial review of the PEM CICEL PROCs were conducted. Comments have
been submitted. The converted STOL procedures are not acceptable for use at
the current time.
PEM is attempting to gain access to the UARS Orbit and Attitude data
files and the OA Services routines. PEM is unable to process without UARS
Orbit and Attitude data.
PEM awaits support of the UARS Traceability Mission for the TIMED mission.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG continues to serve as the point-of-contact
for UARS instrument operations to support re-engineering
of the UARS Mission Operations Concept to develop
capabilities for the Science Traceability Mission.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 05 January 2002 (Orbit 56418, GMT Day 005)
through 11 January 2002 (Orbit 56522, GMT Day 011).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere
viewing) and is performing nominally under two-battery
operations. The instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
---------- ---- ----- -------
HALOE 006/1748 56443 Power OFF
PEM/AXIS 008/0117 56463 AXIS 1 and 2 ON
The beta angle ranged between the angles of 41.18 degrees to
70.28 degrees for this report period. The beta angle is now
decreasing towards a minimum angle of 0 degrees on 03 February
2002 (DOY 034).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3
are currently maintaining greater than 25.60 V end-of-night (EON)
load bus voltage. Battery temperatures are stable with a
temperature delta between Battery 2 and Battery 3 of 0.94 to 0.94
degrees C. Additional battery performance data for each battery
on 22 December 2001 and 28 December 2001 are provided in Appendix
D.
The clock error ranged from -11.1 to 12.0 msec during this report
period. Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the
initial excursion of ALERTS or Out-Of-Limits conditions detected
by the FOT for this report. A summary is listed below and the
details of ALERT and Out-Of-Limits occurrences are listed in
Appendix B.
OUT-OF-LIMITS = 00
ALERTS = 00
The following real-time command sheets were executed for the
listed operational element(s) during this report period (see
Appendix A).
FOT = 5
PEM/AXIS = 1
HALOE = 1
UARS Anomaly Reports initiated during this report period are
listed in Appendix E and summarized below.
FOT = 0
Zero (0) Anomaly Reports remained OPEN at the end of this report
period (see Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 01/333 - 2287.496855
Transponder B: 01/324 - 2287.498688
GROUND SYSTEM OPERATION
Zero (0) operational support problems occurred this period.
Anomaly Reports are included in Appendix E.
The data loss calculations are no longer being supplied by Data
Capture Facility (DCF). This service should resume soon.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
A realtime test was executed on the Eclipse system on 09 January
2002. This test involved sending a single NOOP command from the
Eclipse system to the spacecraft. The command was previously
tested on the simulator and bit busted coming out of the Eclipse
to verify that it worked properly. The command worked as
expected and verified correctly onboard. This was the first
milestone in a string of many to come, in order to get Eclipse
ready for operations.
The FOT ended its testing of the TAC-RS with the legacy AP
system. Through several tests the TAC developers and FOT have
determined that the problem resides in the interface between the
AP and TAC-RS. This interface is currently an RS-422 interface
and will be going away once the Eclipse system goes online.
Therefore, no more daily passes will be taken with the AP and
TAC-RS system. The TAC-RS will be strictly used with the Eclipse
system.
MEETINGS
The FOT attended a UARS Ground System Rehost Meeting on Monday,
January 07, 2002.
The FOT attended a UARS Eclipse Working Meeting on Thursday,
January 10, 2002.
The FOT attended a UARS SSPP Pointing Meeting on Thursday,
January 10, 2002.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 03 February 2002 (DOY
034). This maneuver will be a Reverse to Forward Flight maneuver
(Southern Hemisphere viewing).
Attachments:
Appendix A, Real-time Command Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 11 January 2002
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
56419 005/0254 02-003 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
56443 006/1748 02-004 HALOE Power OFF
56444 006/1929 02-005 OBC/CLK Clk Adj fr 20.6 to 0.00 msec/day
56463 008/0117 02-006 PEM/AXIS AXIS 1 and 2 ON
56473 008/1747 02-007 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
56487 009/1009 02-008 OBC/CDH NOOPS - Test Commands
56511 011/0626 02-009 OBC/CLK Clk Adj 20.6 to 41.2 msec/day
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 11 January 2002
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period.
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period.
_________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 11 January 2002
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
56419 005/0254 Clk Adj 0.00 to 20.6 msec/day
56444 006/1929 Clk Adj fr 20.6 to 0.00 msec/day
56473 008/1747 Clk Adj 0.00 to 20.6 msec/day
56511 011/0626 Clk Adj 20.6 to 41.2 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 11 January 2002
04 January 2002 - GMT Day 004 Beta = 41.18 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ---------- ----- ----- -----
1 -504.0/-728.0 -3.77/-4.69 +0.0/ -0.0 10.72 0.000 0.00
2 +112.0/-952.0 +5.14/+4.48 +26.8/-14.8 25.76 1.030 20.39
3 +224.0/-224.0 +4.48/+3.17 +26.4/-14.8 25.60 1.020 20.39
11 January 2002 - GMT Day 011 Beta = 70.28 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ---------- ---------- ----- ----- -----
1 -504.0/-672.0 -3.16/-3.77 +0.0/ -0.0 10.72 0.000 0.00
2 +000.0/-224.0 +2.85/+2.52 +17.6/-12.8 27.36 1.030 14.33
3 +224.0/-224.0 +2.20/+1.55 +16.8/-12.0 27.36 1.030 13.98
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 11 January 2002
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period.
SEVNTFW02/02